OpenStudy (anonymous):

hey I'm looking for a solution of the ordinary differential equation y'' -c*y'+10=0 with y(0)=a and y(T)=b. Can anyone help me out? Thanks!

OpenStudy (anonymous):

dy/dx=p hence dp/(c*p+10)=dx then u have: ln(c*p+10)=c*x+A A is a constant or c*p+10=exp(c*x+A) c*(dy/dx)+10=exp(c*x+A) dy/dx=(exp(c*x+A)-10)/c* the final solution is: y=(1/c*^2)exp(c*x+A)-(10/c*)x+B B is a constant now you can use boundery conditoins

OpenStudy (anonymous):

y'' -c*y'+10=0 or y'' -c*y'+10y=0

OpenStudy (anonymous):

the problem is as I stated it, y'' -c*y' +10 = 0 i.e. it's inhomogeneous I believe. But your answer helps a lot! I think I can take it from there.

OpenStudy (anonymous):

ok